OpenStudy (anonymous):

Forty-five is 30% of what number? A. 45 B. 150 C. 67 D. 135

OpenStudy (anonymous):

\[45 = \frac{ 30 }{ 100} n\]

OpenStudy (anonymous):

What did you get, Koi?

OpenStudy (anonymous):

I got 135

OpenStudy (anonymous):

so what option is it?

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

45/30 =3/2 so the number will be (3/2)*100=300/2=150 is it correct

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

wait.....no...

OpenStudy (anonymous):

n should be 150

OpenStudy (anonymous):

100*45/30=150 not 135

OpenStudy (anonymous):

oh yea....

OpenStudy (anonymous):

^

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

what's wrong with my calculation lately...

OpenStudy (anonymous):

Koi, it's not 135 it's 150! hani deserves a medal!

OpenStudy (anonymous):

thanks